OpenStudy (anonymous):

Find the radius of convergence and the interval of convergence of the series. Series in comments

OpenStudy (anonymous):

\[\sum_{k=1}^{\infty} \frac{ x^k }{ 2\times4\times6...(2k) }\]

OpenStudy (anonymous):

looks like a good candidate for the ratio test

OpenStudy (anonymous):

\[\frac{a_{n+1}}{{a_n}}=\frac{x}{2k+2}\] i think

OpenStudy (anonymous):

did you just plug in k+1 and then cancel out with k?

OpenStudy (amistre64):

essentially \[\frac{ \cancel{2\times4\times6...(2k) }}{ \cancel{2\times4\times6...(2k)}~(2(k+1)) }=\frac{1}{2k+2}\]

OpenStudy (anonymous):

So I would just find the limit after that?

OpenStudy (amistre64):

yep, which can then be assessed for radius and interval convergences

OpenStudy (amistre64):

|x|*0 <1 .... always so the radius of convergence is: infinity, and the interval is (-inf,inf)